solve the equation; 3x-4-5x=x+4+x. i understand the basic principle of how to work the problem out but not understanding how to get the answer given.
Answer by rfer(16322)

-2x-4=2x+4
-2x-2x=4+4
-4x=8
x=8/-4
x=-2